NicholasH:

12th Grade Math, (Mathematics for Data and Financial Literacy) Evaluate the numerical expression the quantity 5 to the power of five sixths end quantity over the quantity 5 to the power of one sixth end quantity. square root of 15 square root of 125 cube root of 10 cube root of 25

NicholasH:

jhonyy9:

how you know this rule of powers when you divide the same terms like in your case 5 on different powers this is the formula \[\frac{ x^{a} }{ x^{b} } = x^{(a-b)}\] have you learn about this ?

jhonyy9:

so to get the correct answer you need use this rule first

NicholasH:

There is a lot of formulas like this, what do i do with this?

jhonyy9:

in ths first step you need calcule \[\frac{ 5 }{ 6 } - \frac{ 1 }{ 6 } = ?\]

NicholasH:

how you know to set it up like that?

NicholasH:

6/6?

NicholasH:

ok

NicholasH:

?

jhonyy9:

@nicholash wrote:
6/6?
how you get 6/6 ?

NicholasH:

i meant 4/6

jhonyy9:

yes so is right

NicholasH:

ok, so what after that then?

jhonyy9:

but 4/6 you can simplify it

NicholasH:

yeah to 2/3

jhonyy9:

exactly

NicholasH:

so how do we get it to square root?

jhonyy9:

so will get 5 on power of 2/3 about what i ve said that the denominator of a fraction power is the indice of radical so in this case how can you rewrite your result ?

NicholasH:

3 square root 5?

NicholasH:

meant 10

jhonyy9:

no 5 will be on power of 2 what is equal 5*5 = ? and the index of radical is 3 what mean cube root

NicholasH:

oh so 3 with 25 in the box?

jhonyy9:

this 3 mean cubroot and yes 5 squared is 25

jhonyy9:

from these choices what is right ?

NicholasH:

so it is d?

jhonyy9:

perfect do you understand it now sure ?

NicholasH:

I have a understanding of it anyway

jhonyy9:

ok cya

NicholasH:

thank you.